P(x): x2/2 = x

Find the following then identify their truth values.

  1. P (1)
  2. P (2)
  3. An, P(n)
  4. En, P(n)




1
Expert's answer
2022-04-06T09:10:24-0400

First of all, solve the equation: "\frac{x^2}{2}=x". The latter can be written as: "(\frac12x-1)x=0". Solutions of the latter are "x=2" and "x=0." Thus, we receive the following:

  1. "P(1)" is false. 2. "P(2)" is true. 3. "\forall n\,\, P(n)" is false (not all numbers\natural numbers are solutions). 4. "\exists n\,\,P(n)" is true, since "x=2" is a solution.